OpenStudy (anonymous):

Please help me, will give medal. y = 6 – x and y = x – 2 (1 point) (2, 4) (–4, 2) (4, 2) no solutions

OpenStudy (vishweshshrimali5):

You only have to solve the 2 equations. Best method is just to add both equations and obtain the value of y and then with the help of any one equation and the afore mentioned value of y obtain the value of x.

OpenStudy (anonymous):

so it is c, right? Also could you help me with some other questions?

OpenStudy (vishweshshrimali5):

You can post your question so that we can help you out.

OpenStudy (anonymous):

y = 2x – 1 and y = x + 3 (1 point) (4, 7) (7, 4) (–7, –4) infinite solutions

OpenStudy (vishweshshrimali5):

First show me what your method is.

OpenStudy (anonymous):

you can do many different things like just plug in the numbers.. what is it that you need exactly (7) = 2(4)

OpenStudy (anonymous):

Oh sorry posted the wrong question, already got that one right. What will the graph look like for a system of equations that has no solution? (1 point) The lines will be perpendicular. The lines will cross at one point. Both equations will form the same line. The lines will be parallel.

OpenStudy (anonymous):

I think it is d

OpenStudy (vishweshshrimali5):

Yes u r correct.

OpenStudy (anonymous):

This is one question that i really need help with. What is the rule for the function shown in the table? (1 point) x = -1, 0, 1, 2 y = -2, 1, 4, 7 A.y = 1/3x + 1 B.y = 3x + 1 C.y = 1/3x D.y = 1/2x + 1